The sum of a number cubed and 3 times the square of the number is 75 more than 25 times the number. Find the number. HInt: Let X= the number
Answer by mananth(16949)

x^3+3x^2=25x+75
x^3+3x^2-25x-75
>>
when you substitute x = -5 the expression is =0
(x+5) is a root
(x+5)(x-5)(x+3)=0
...
5 , -5 ,-3 are the roots of the equation.
check by plugging the value in the equation to find the number
..
m.ananth@hotmail.ca